# MACRO-LLM: LLM-Empowered Multi-Agent Collaborative Reasoning under Spatiotemporal Partial Observability

### v1

@@ 3.4 Introspector
Due to spatiotemporal partial observability, an agent’s state inference is often compromised by dynamic environmental shifts and the unpredictable behaviors of others. Consequently, if an agent observes a reward decline between time steps, it initiates a self-reflection mechanism to analyze whether substantial environmental changes have occurred. We model this process as a Semantic Gradient Descent, utilizing the reward gap as a signal to update spatiotemporal strategies.
Observation-Consistency-based Adaptive Learning Rate.
We first quantify the magnitude of environmental shifts to determine the "step size" for strategy updates. Let $o_{n}^{t}$ denote the vectorized observation of agent $n$ at time $t$. We construct a transition vector $\Gamma_{n}^{t}$ by concatenating the observation difference and the action vector, defined as $\Gamma_{n}^{t}=(\Delta{o_{n}^{t}},a_{n}^{t})$, where $\Delta{o_{n}^{t}}=o_{n}^{t+1}-o_{n}^{t}$.
To evaluate scene consistency, we compute the cosine similarity between two consecutive transitions, $\Gamma_{n}^{t-1}$ and $\Gamma_{n}^{t}$. The reflection learning rate ($lr$) is derived from scene discontinuity:
$lr=1-\frac{\Gamma_{n}^{t-1}\cdot\Gamma_{n}^{t}}{\|\Gamma_{n}^{t-1}\|\cdot\|\Gamma_{n}^{t}\|}.$ (5)
Here, $lr\in[0,2]$ is a scalar inversely proportional to transition similarity. A higher $lr$ indicates a drastic environmental shifts, requiring aggressive strategy updates, whereas a lower $lr$ implies minor fluctuations, requiring only fine-grained refinements.
Semantic Gradient Computation.
Inspired by textual gradient descent Yuksekgonul et al. (2025), we formalize the semantic gradient as a natural language description of the "direction" for strategic improvement. By analyzing observation logs and negotiation records $\mathbf{N}$ (where $\mathbf{N}_{t}$ denotes the negotiation history at time $t$), the agent identifies the root cause of the performance drop. The semantic gradient is generated via: $g_{n,t}^{\text{semantic}}=\text{LLM}\left(\Gamma_{n}^{t-1},\Gamma_{n}^{t},\Delta R_{n}^{t},\Pi_{n}^{\text{temp}},\Pi_{n}^{\text{spatial}},\mathbf{N}_{t-1},\mathbf{N}_{t}\right)$. Here, $g_{n,t}^{\text{semantic}}$ represents textual rules that explain the causal links between actions, state changes, and negative reward outcomes, effectively pointing out the flaws in the current strategy.
@@
Finally, we update the temporal strategy using the computed gradient and learning rate. The update rule is conceptually formulated as:
$\Pi_{n,t+1}^{\text{temp}}\leftarrow\Pi_{n,t}^{\text{temp}}-lr\cdot g_{n,t}^{\text{semantic}}.$ (6)
Equation (6) represents a prompt-based operation where the scalar $lr$ modulates the instruction intensity: a high $lr$ prompts the LLM to "discard and rewrite" the strategy based on $g_{n,t}^{\text{semantic}}$, while a low $lr$ implies a "fine-grained refinement" of the existing $\Pi_{n,t}^{\text{temp}}$. This process completes the semantic learning loop, enabling MACRO-LLM to adapt to dynamic environments by emulating stochastic gradient descent within the semantic strategy space.
@@ 4 Experiments
We evaluate MACRO-LLM on two distinct interactive tasks requiring complex multi-turn coordination under partial observability: Cooperative Adaptive Cruise Control (CACC) Chu et al. (2020) and Pandemic Control (PC) Kompella* et al. (2020).
@@ 4.1 Experimental Settings
Table 1: Performance comparison on the CACC task across Catch-up and Slow-down scenarios. Metrics include RMSE and SD for Headway (H) and Velocity (V). Bold and underline indicate best and second-best results, respectively.
@@
We implement the MACRO-LLM framework using GPT-4o (gpt-4o-2024-08-06) via the OpenAI API with the openai Python library (version 1.75.0). To ensure focused reasoning and reproducibility, we set the sampling temperature to 0.3 and top-p to 1.0. All simulation environments and local baselines (e.g., MARL algorithms) were executed on a server equipped with 4 NVIDIA GeForce RTX 3090 GPUs. Detailed prompt templates are provided in Appendix J.
Table 2: Performance comparison on the PC task across Helsinki, Hong Kong, and New York topologies. Metrics include Normalized Infection (I_n), Peak Infection (PI_n), Normalized Dead (D_n), and Pandemic Duration (PD). Bold and underline indicate best and second-best results, respectively.
@@
Task Configurations.
We evaluate MACRO-LLM across two diverse domains requiring complex coordination. (1) CACC: A continuous vehicle platooning task where a fleet of agents (default $N=8$) must maintain safety and efficiency during dynamic "Catch-up" and "Slow-down" scenarios under partial observability. (2) Pandemic Control (PC): A strategic planning task where agents (buildings) regulate mobility to mitigate virus spread across three urban topologies (Helsinki, Hong Kong, and New York). These tasks evaluate the framework’s capability in both fine-grained control and long-horizon strategic planning. Refer to Appendix D for detailed simulator configurations and city topologies.
@@
We benchmark MACRO-LLM against representative methods from two paradigms: (1) LLM-based MAS baselines: We select representative frameworks including ToM-Belief Li et al. (2023b), ChatEval Chan et al. (2023), and LAMEN Davidson et al. (2024); and (2) MARL baselines: We select representative MARL algorithms including DPPO, DMPO Ma et al. (2024), and IC3Net Singh et al. (2018). Details on baseline implementations, prompt adaptations, and fair comparison settings are provided in Appendix E.
@@
For the CACC task, we evaluate platoon stability via the Root Mean Square Error (RMSE) and Standard Deviation (SD) of both velocity and headway distance (denoted as RMSE-V/H and SD-V/H). For the PC task, we assess containment efficacy through Normalized Infection (I_n), Peak Infection (PI_n), Normalized Death (D_n) and total Pandemic Duration (PD). Metric definitions are detailed in Appendix F.
Table 3: Generalizability analysis of MACRO-LLM across different LLM backbones. S1/S2: CACC Catch-up/Slow-down scenarios. C1-C3: PC topologies for Helsinki, Hong Kong, and New York.
Metrics ($\downarrow$)
Tasks Base Models RMSE-H RMSE-V SD-H SD-V
CACC S1 DeepSeek-V3.1 1.528 0.553 0.611 0.190
Qwen3-flash 3.521 1.910 2.935 0.579
GPT-4o-mini 2.906 1.667 1.858 0.936
GPT-4o 1.212 0.897 0.455 0.283
S2 DeepSeek-V3.1 0.649 3.388 0.274 0.417
Qwen3-flash 3.871 3.678 2.946 0.974
GPT-4o-mini 2.352 3.999 1.598 1.171
GPT-4o 0.561 3.325 0.209 0.345
Tasks Base Models I_n PI_n D_n PD
PC C1 DeepSeek-V3.1 0.098 0.038 0.002 46
Qwen3-flash 0.010 0.008 0.002 16
GPT-4o-mini 0.010 0.010 0.000 17
GPT-4o 0.010 0.008 0.000 9
C2 DeepSeek-V3.1 0.008 0.005 0.001 22
Qwen3-flash 0.008 0.006 0.001 18
GPT-4o-mini 0.006 0.004 0.000 24
GPT-4o 0.006 0.005 0.001 22
C3 DeepSeek-V3.1 0.003 0.003 0.000 11
Qwen3-flash 0.003 0.003 0.000 13
GPT-4o-mini 0.003 0.003 0.000 13
GPT-4o 0.005 0.003 0.000 33
4.2 Performance of CACC Tasks
Based on the experimental results presented in Table 1, MACRO-LLM demonstrates superior coordination capabilities, particularly in enhancing platoon stability and consistency compared to both MARL-based and LLM-based baselines. (1) In the Catch-up scenario, MACRO-LLM reduces the SD-H by 86.3% compared to LAMEN and 54.2% compared to DPPO. Furthermore, it significantly improves tracking precision, reducing RMSE-H by 24.4% relative to DPPO. This indicates that the negotiation mechanism of MACRO-LLM effectively mitigates the oscillations observed in baselines, resulting in smoother and safer acceleration phases. (2) In the Slow-down scenario, the results highlight a strategic trade-off where MACRO-LLM prioritizes headway maintenance for safety over aggressive velocity matching. While maintaining competitive velocity errors, our framework substantially outperforms baselines in maintaining safe distances. MACRO-LLM improves SD-V by 28.1% over the ToM-Belief. These results collectively validate that MACRO-LLM ensures robust performance.
Figure 3: Scalability analysis of MACRO-LLM on the CACC task for number of agents $N\in\{8,16,24,32\}$.
4.3 Performance on PC Tasks.
Table 2 presents the comparative results across three distinct city topologies: Helsinki, Hong Kong, and New York. MACRO-LLM consistently demonstrates superior containment efficacy, achieving the lowest normalized infection rates (I_n) and peak infections (PI_n) while significantly shortening the pandemic duration (PD) across all settings. Specifically, in the most complex New York scenario, MACRO-LLM significantly outperforms DPPO, reducing cumulative infections by over 99.2% and shortening the pandemic duration by nearly 3 timestimes compared to DPPO. This indicates that our framework effectively balances peak suppression with rapid containment across varying population scales. Notably, while baselines often struggle with the trade-off between strict containment and duration, MACRO-LLM optimizes both simultaneously. For instance, in Hong Kong, while LAMEN achieves low infection rates, MACRO-LLM further reduces the pandemic duration, demonstrating more efficient long-term planning. This robust performance highlights the critical role of the Negotiator and Introspector modules in handling complex, large-scale coordination.
4.4 Analysis of Model Generalizability
We investigate the robustness of MACRO-LLM by instantiating it with diverse LLM backbones, including GPT-4o-mini, Qwen3-flash Bai et al. (2023), GPT-4o, and DeepSeek-V3.1 Liu et al. (2024). The results are shown in Table 3. For the precision-critical CACC task, fine-grained spatial control is heavily influenced by the reasoning capability of the base model. Powerful models like GPT-4o and DeepSeek-V3.1 significantly reduce RMSE-H by over 47.4% in the Catch-up scenario (S1) compared to GPT-4o-mini. Conversely, in the PC task, Qwen3-flash achieves containment efficacy comparable to GPT-4o. This demonstrates that MACRO-LLM effectively compensates for the limitations of smaller models in strategic planning tasks, demonstrating the framework’s robustness and feasibility across various computational budgets.
@@ 4.5 Analysis of Scalability
To investigate the scalability of MACRO-LLM, we extended the CACC Catch-up scenario to varying fleet sizes of $N\in\{8,16,24,32\}$, analyzing the headway and velocity trajectories of representative agents located at the front, middle, and rear of the platoon. As illustrated in Figure 3, MACRO-LLM maintains consistent robustness to increasing fleet sizes, even as the network quadruples to $N=32$. While the propagation delay naturally increases with platoon size, the oscillation amplitudes remain bounded and effectively dampened over time. All agents, regardless of their position, converge to the target equilibrium within 40s. By compressing the dense interactions of a large-scale network into aggregated statistical features, each agent only needs to process a fixed-size local context and a virtual mean-field node, thereby mitigating communication overhead and contextual overload. The results demonstrate that MACRO-LLM effectively balances local precision with global coherence, ensuring scalability for large-scale deployment. More analysis about scalability and cost (including communication and computation cost) are provided in Appendix G.
@@ 4.6 Ablation Study
Figure 4: Ablation of MACRO-LLM—impact of each module on CACC performance.
Figure 5: Ablation of MACRO-LLM—impact of each module on PC performance
We validate the contribution of each module in Figs. 4 and 5. As shown in Fig. 4, removing the CoProposer leads to immediate coordination failure, resulting in collisions in CACC. Ablating the Negotiator causes stability metrics to degrade drastically, demonstrating that unstructured communication fails to resolve conflicting intentions. Furthermore, without the Introspector, agents struggle to suppress oscillations over time due to the lack of self-reflection. In the PC task (Fig. 5), this manifests as longer PD and higher $I_{n}$. Comprehensive quantitative analyses are provided in Appendix H.
@@ 5 Conclusion
In this work, we present MACRO-LLM, a novel framework enabling agents to reason and coordinate through localized negotiations under spatiotemporal partial observability. We explicitly decompose the coordination challenge along spatial and temporal dimensions, addressing them through three interconnected modules: (1) a CoProposer to generate collaborative actions and mitigate future uncertainty via predictive rollouts; (2) a Negotiator that resolves spatial conflicts using mean-field statistical estimations; and (3) an Introspector that facilitates strategic refinement via semantic gradient descent. Extensive experiments on CACC and PC tasks demonstrate MACRO-LLM’s capability to align local actions with long-term global objectives, achieving superior collaborative efficiency compared to representative baselines. Ultimately, our work highlights the critical role of explicitly modeling spatiotemporal dynamics in scaling LLM-based multi-agent systems.
@@
While MACRO-LLM demonstrates robust coordination capabilities, we acknowledge two primary limitations. First, the iterative negotiation mechanism introduces computational overhead and latency, necessitating a trade-off between reasoning depth and inference speed for time-sensitive applications. The theoretical communication and computation cost analyses are provided in Appendix G. Second, relying on foundation model APIs requires additional context injection for domain adaptation, which increases token consumption and inference costs. Future work will explore lightweight fine-tuning or Retrieval-Augmented Generation (RAG) to internalize domain knowledge, thereby enhancing execution efficiency and reducing context dependency.
@@ Appendix A Extended Related Works
There has been a recent surge in research on the reasoning, planning, and decision-making capabilities of LLMs to construct agents across diverse domains Huang et al. (2024). Early works, such as Chain-of-Thought (CoT) Wei et al. (2022) and ReAct Yao et al. (2023), enhance reasoning by decomposing complex tasks into subtasks. Another common approach involves generating a broader set of candidate plans through multiple reasoning steps to widen the exploration of possible solutions such as Self-Consistency Wang et al. (2022b) and Graph-of-Thought Besta et al. (2024). To further strengthen reasoning, Self-Refine Madaan et al. (2023) applies self-reflection module, prompting agents to critique failures and iteratively update strategies. To enhance reasoning with historical knowledge, MemoryBank Zhong et al. (2024) leverages memory-augmented planning to retrieve relevant knowledge and past experiences for guiding decision-making. These diverse methods can be integrated into one agent pipeline to collectively enhance overall capacity. Nevertheless, integrating multiple sophisticated components often incurs high computational costs and can be constrained by context length limitations, potentially undermining coherence in long-horizon tasks.
@@
(1) Hierarchical aggregation: Works such as FinCon Yu et al. (2024), MLPO Estornell et al. (2025), and HALO Hou et al. (2025) employ hierarchical architectures. These systems designate specific agents (e.g., central hubs) to aggregate incomplete information, constructing a pseudo-global view to compensate for limited local perception. However, this centralization introduces critical bottlenecks: it relies heavily on central nodes and often incurs prohibitive communication overheads as the network scales. (2) Individual reasoning enhancement: A complementary line of research focuses on enhancing individual reasoning capabilities. Theory-of-Mind (ToM) approaches Li et al. (2023b) enable agents to infer the beliefs and intents of others, while frameworks like LAMEN Davidson et al. (2024) utilize structured reasoning chains to predict hidden states. While effective for small-scale interactions, these methods often struggle with the spatiotemporal uncertainty inherent in complex topologies, where agents must simultaneously reason about spatial neighbors and future dynamics under strict latency constraints.
@@
To improve sample efficiency, advanced policy optimization methods such as Multi-Agent PPO (MAPPO) Schulman et al. (2017) and Model-based Decentralized Policy Optimization (DMPO) Ma et al. (2024) have been developed. Furthermore, adaptive communication protocols like IC3Net Singh et al. (2018) introduce gating mechanisms to decide when to communicate. Limitations: Despite these advances, the fundamental reliance on trial-and-error interaction makes MARL computationally intensive (see Fig. 7). As noted by Huh and Mohapatra (2023) and Oroojlooy and Hajinezhad (2023), MARL policies often struggle to generalize to unseen topologies or agent numbers, necessitating the retraining costs that our LLM-based framework aims to eliminate.

### v2

@@ 3.4 Introspector
Due to spatiotemporal partial observability, an agent’s state inference is often compromised by dynamic environmental shifts and the unpredictable behaviors of others. Consequently, if an agent observes a reward decline between time steps, it initiates a self-reflection mechanism to update strategy. We model this process as scenario-aware adaptive reflection, which determines how much to revise based on environmental drift and what to revise through attribution-based reasoning.
Revision Intensity Estimation.
We first quantify the magnitude of environmental shifts to determine the revision intensity. Let $o_{n}^{t}$ denote the vectorized observation of agent $n$ at time $t$. We construct a transition vector $\Gamma_{n}^{t}$ by concatenating the min-max normalized observation difference and the action vector, defined as $\Gamma_{n}^{t}=(\Delta{\tilde{o}_{n}^{t}},\tilde{a}_{n}^{t})$, where $\Delta{o_{n}^{t}}=o_{n}^{t+1}-o_{n}^{t}$.
To evaluate scene consistency, we compute the cosine similarity between two consecutive transitions, $\Gamma_{n}^{t-1}$ and $\Gamma_{n}^{t}$. The drift intensity ($\lambda_{\text{drift}}$) is derived from scene discontinuity:
$\lambda_{\text{drift}}=1-\frac{\Gamma_{n}^{t-1}\cdot\Gamma_{n}^{t}}{\|\Gamma_{n}^{t-1}\|\cdot\|\Gamma_{n}^{t}\|}.$ (5)
Higher $\lambda_{\text{drift}}$ indicates greater environmental drift requiring aggressive revision, while lower values permit only fine-grained refinements. Combined with the reward-decline-based trigger condition to ensure genuine performance degradation.
Revision Direction Identification.
We generate a semantic revision signal as a natural language description of the “direction” for strategic improvement. The LLM analyzes observation logs and negotiation records $\mathbf{N}_{t}$ to identify which transitions deviated from the temporal strategy and to assess the impacts from neighbors. The revision signal $g_{n,t}^{\text{semantic}}$ is generated as textual rules that characterize the associations between actions, state changes, and negative reward outcomes, providing diagnostic guidance for strategy refinement.
@@
Finally, we update the temporal strategy using the revision signal and drift intensity. A concrete prompt template for strategy update with revision intensity and direction is provided in Appendix Q.3. The update is formulated as $\Pi_{n,t+1}^{\text{temp}}\leftarrow\text{LLM}\left(\Pi_{n,t}^{\text{temp}},\ g_{n,t}^{\text{semantic}},\ \lambda_{\text{drift}}\right)$. The drift intensity $\lambda_{\text{drift}}$ modulates the revision scope: a high $\lambda_{\text{drift}}$ indicates substantial environmental drift, instructing the LLM to aggressively revise the strategy according to $g_{n,t}^{\text{semantic}}$; a low $\lambda_{\text{drift}}$ confines revision to fine-grained adjustments, preserving the core of the existing strategy $\Pi_{n,t}^{\text{temp}}$.
@@ 4.1 Experimental Settings
Table 1: Performance comparison on the CPP task across Catch-up and Slow-down scenarios. Metrics include RMSE and SD for Headway (H) and Velocity (V). Bold and underline indicate best and second-best results, respectively.
@@
[…]
MACRO-X: DS denotes DeepSeek-V3.1, Llama denotes
Llama-3.3-70B-Instruct, and GPT denotes GPT-4o.
Task Configurations.
We evaluate MACRO-LLM across two complementary tasks where coordination outcomes are measured through domain-specific physical metrics, ensuring that performance gains reflect genuine coordination ability rather than language fluency. (1) Cooperative Platoon Planning (CPP) Chu et al. (2020): a multi-agent vehicle platooning task (default N=8) with dynamic “Catch-up” and “Slow-down” scenarios. (2) Pandemic Control (PC) Kompella* et al. (2020): a strategic planning task where agents select regulation level to mitigate virus spread across three urban topologies (Helsinki, Hong Kong, and New York). CPP and PC tasks span continuous and discrete action spaces, linear and mixed topologies, and fine-grained and long-horizon decision-making, respectively. Refer to Appendix G for simulator configurations and city topologies.
@@
We implement MACRO-LLM as a backbone-agnostic framework and evaluate it across six LLM backbones: GPT-4o, GPT-4o-mini, DeepSeek-V3.1, Llama-3.3-70B-Instruct, Mistral-Small-3.2-24B-Instruct, and Qwen3-flash. Tables 1 and 2 report three representative backbones; the full generalizability analysis is in Appendix K. Scalability, ablation, and robustness analyses use GPT-4o as the default. Framework and task configurations are detailed in Appendix G. Results for MACRO-LLM are averaged over 5 independent runs; variance analysis confirming low standard deviations across all metrics is provided in Appendix N. A case study on hierarchical deployment with different LLM invocation frequency is in Appendix I.
Table 2: Performance comparison on the PC task across Helsinki, Hong Kong, and New York topologies. Metrics include Normalized Infection (I_n), Peak Infection (PI_n), Normalized Death (D_n), and Pandemic Duration (PD). Bold and underline indicate best and second-best results, respectively.
@@
[…]
MACRO-X: DS denotes DeepSeek-V3.1, Llama denotes Llama-3.3-70B-Instruct, and GPT denotes GPT-4o.
@@
We benchmark MACRO-LLM against representative methods from two paradigms: (1) LLM-based MAS baselines, which share the same backbone (GPT-4o) and environmental interfaces to provide controlled comparisons: ToM-Belief Li et al. (2023b), ChatEval Chan et al. (2023), and LAMEN Davidson et al. (2024); and (2) MARL baselines, which serve as task-specialized performance references trained on each topology: DPPO, DMPO Ma et al. (2024), and IC3Net Singh et al. (2018). Details on baseline implementations, prompt adaptations, and the division between LLM reasoning and external numerical modules are provided in Appendix H.3.
@@
For the CPP task, we evaluate platoon stability via the Root Mean Square Error (RMSE) and Standard Deviation (SD) of both velocity and headway distance (denoted as RMSE-V/H and SD-V/H). For the PC task, we assess containment efficacy through Normalized Infection (I_n), Peak Infection (PI_n), Normalized Death (D_n) and total Pandemic Duration (PD). Metric definitions are detailed in Appendix J.
4.2 Performance of Cooperative Platoon Planning Task
As shown in Table 1, MACRO-LLM consistently outperforms both MARL and LLM-based baselines across multiple backbones. In the Catch-up scenario, all three MACRO-LLM variants achieve the top RMSE-H and SD-H results. The SD-H of MACRO-GPT ($0.409\pm 0.034$, Appendix N) is 87.7% lower than that of LAMEN and 58.9% lower than that of DPPO, while even MACRO-DS attains lower SD-H than all baselines. In the Slow-down scenario, MACRO-LLM prioritizes headway safety over aggressive velocity matching. MACRO-GPT achieves the lowest SD-H ($0.128\pm 0.054$, Appendix N), a 43.1% improvement over ToM-Belief, while maintaining competitive velocity errors. The consistent advantage across three diverse backbones confirms that the performance gains stem from the framework’s negotiation mechanism rather than any single model’s capability.
Figure 3: Scalability analysis of MACRO-LLM on the CPP task for number of agents $N\in\{8,16,24,32\}$.
4.3 Performance on Pandemic Control Task.
Table 2 presents the comparative results across three city topologies. MACRO-LLM consistently achieves the lowest infection rates and shortest pandemic durations across all settings. In New York topology, all three MACRO-LLM variants reduce cumulative infections by over 99% compared to DPPO while shortening the pandemic duration from 95 to as few as 11 days (MACRO-DS). In Hong Kong, MACRO-Llama reduces PD to 16 days, outperforming LAMEN (23 days) at comparable infection levels. MACRO-LLM achieves both low infections and short durations simultaneously. This consistent advantage across three diverse backbones confirms that the coordination gains stem from the framework’s negotiation mechanism rather than any single model’s capability.
@@ 4.4 Analysis of Scalability
To investigate scalability, we extend the CPP Catch-up scenario to fleet sizes $N\in\{8,16,24,32\}$. As illustrated in Fig. 3, MACRO-LLM maintains consistent robustness as the network quadruples. Oscillation amplitudes remain bounded and all agents converge to the target equilibrium within 40s regardless of fleet size. Aggregated metrics (Appendix L, Table 6) reveal two complementary trends: RMSE-H decreases from 1.237 (NN=8) to 0.728 (NN=32), as the growing proportion of already-stabilized agents dilutes the transient deviations of those still responding to propagation. Meanwhile, SD-H increases moderately from 0.409 to 0.433, reflecting the inherently longer synchronization lag in larger platoons. Importantly, this increase is bounded. At NN=32, SD-H remains 56.4% lower than that of DPPO (NN=8), confirming that MACRO-LLM compresses large-scale interactions while ensuring coordination quality. Further scalability analyses are provided in Appendix L.
@@ 4.5 Ablation Study
Figure 4: Ablation of MACRO-LLM—impact of each module on CPP performance.
Figure 5: Ablation of MACRO-LLM—impact of each module on PC performance.
We validate each module’s contribution in Figs. 4 and 5. Ablating the Negotiator degrades stability metrics drastically. Without the Introspector, agents fail to suppress oscillations over time. Notably, the ablations reveal cross-module dependencies: removing the CoProposer disables not only temporal verification but also the Negotiator’s confidence assessment without rollout rewards. It explains the largest SD-H degradation among all variants. Quantitative analyses are in Appendix M.
@@ 5 Conclusion
In this work, we present MACRO-LLM, a novel framework enabling agents to reason and coordinate through localized negotiations under spatiotemporal partial observability. We explicitly address the coordination challenge along spatial and temporal dimensions, addressing them through three interdependent modules: (1) a CoProposer to generate collaborative actions and mitigate future uncertainty via predictive rollouts; (2) a Negotiator that resolves spatial conflicts using mean-field statistical estimations; and (3) an Introspector that facilitates strategic refinement via scenario-aware adaptive reflection. Extensive experiments on CPP and PC tasks demonstrate MACRO-LLM’s capability to align local actions with long-term global objectives, achieving superior collaborative efficiency compared to representative baselines. Ultimately, our work highlights the critical role of explicitly modeling spatiotemporal dynamics in scaling LLM-based MAS.
@@
While MACRO-LLM demonstrates robust coordination capabilities, we acknowledge two primary limitations. First, as a strategic-level reasoning framework, MACRO-LLM’s iterative negotiation mechanism prioritizes decision quality over inference speed. The per-step latency is dominated by API network overhead of remote LLM calls, which is orthogonal to the framework’s reasoning complexity and can be mitigated through two composable strategies already validated in our experiments: (1) hierarchical deployment, where MACRO-LLM issues strategic directives every $\Delta$ steps while a lightweight controller maintains millisecond-level safety guarantees. Empirical results confirm that even with 90% fewer LLM invocations, coordination quality surpasses all baselines (Appendix I, Table 4); and (2) local deployment of lightweight models such as Qwen-flash and Mistral-small-3.2-24B-Instruct, which eliminates API latency entirely while maintaining competitive performance (Appendix K). Future work will explore tighter co-design between the strategic layer and domain-specific execution layers to further optimize the reasoning-latency balance. Second, relying on foundation models requires additional context injection for domain adaptation, which increases token consumption. Future work will explore online self-evolution mechanisms, such as experience-driven prompt distillation and adaptive strategy caching, that enable agents to progressively internalize domain-specific knowledge through interaction, reducing context dependency while improving cross-domain adaptability.
Appendix Table of Contents
@@ Appendix A Extended Related Works
There has been a recent surge in research on LLM-based agent reasoning and planning Huang et al. (2024); Yuksekgonul et al. (2025). Techniques such as Chain-of-Thought (CoT) Wei et al. (2022), ReAct Yao et al. (2023), Self-Consistency Wang et al. (2022b), and Graph-of-Thought Besta et al. (2024) enhance single-agent reasoning by decomposing tasks or broadening candidate exploration. Self-Refine Madaan et al. (2023) introduces iterative self-correction, and MemoryBank Zhong et al. (2024) augments planning with historical knowledge retrieval. However, these methods are designed for single-agent settings with full or near-complete observability. In multi-agent scenarios under partial observability, the core challenge shifts from individual reasoning quality to inter-agent coordination under fragmented information. That is a problem that single-agent planning enhancements do not address.
@@
  1. 1.
    Hierarchical aggregation: Works such as FinCon Yu et al. (2024), MLPO Estornell et al. (2025), and HALO Hou et al. (2025) employ hierarchical architectures. These systems designate specific agents (e.g., central hubs) to aggregate incomplete information, constructing a pseudo-global view to compensate for limited local perception. However, this centralization introduces critical bottlenecks: it relies heavily on central nodes and often incurs prohibitive communication overheads as the network scales.
  2. 2.
    Individual reasoning enhancement: A complementary line of research focuses on enhancing individual reasoning capabilities. Theory-of-Mind (ToM) approaches Li et al. (2023b) enable agents to infer the beliefs and intents of others, while frameworks like LAMEN Davidson et al. (2024) utilize structured reasoning chains to predict hidden states. While effective for small-scale interactions, these methods often struggle with the spatiotemporal uncertainty inherent in complex topologies, where agents must simultaneously reason about spatial neighbors and future dynamics under strict latency constraints.
  3. 3.
    Swarm-based optimization: Recent works apply particle swarm optimization to LLM systems. Model Swarms Feng et al. (2024) searches the model weight space over 50 iterations with per-particle validation-set evaluation to produce adapted expert checkpoints. SwarmAgentic Zhang et al. (2025) searches the system configuration space over 10 iterations, jointly optimizing agent functionalities and collaboration structures through failure-driven refinement. While both methods achieve dynamic configuration evolution, this adaptation occurs during an offline optimization loop that requires a globally accessible fitness function evaluated over the entire validation set.
@@
To improve sample efficiency, advanced policy optimization methods such as Multi-Agent PPO (MAPPO) Schulman et al. (2017) and Model-based Decentralized Policy Optimization (DMPO) Ma et al. (2024) have been developed. Furthermore, adaptive communication protocols like IC3Net Singh et al. (2018) introduce gating mechanisms to decide when to communicate. Despite these advances, the fundamental reliance on trial-and-error interaction makes MARL computationally intensive. As noted by Huh and Mohapatra (2023) and Oroojlooy and Hajinezhad (2023), MARL policies often struggle to generalize to unseen topologies or agent numbers, necessitating the retraining costs that our LLM-based framework aims to eliminate.